OpenStudy (anonymous):

a cyclist rides his bike at a speed of 18 miles per hour. what is this speed in kilometers per hour? how many kilometers will the cyclist travel in 4 hours? in your computations, assume that 1 mile is equal to 1.6 kilometer. do not round your answers Speed = km/h distance traveled in 4 hours = km

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

is it 18(1.6) = 28.8 is the speed? and 4(28.8) =115.2 is the distance in 4 hours?

OpenStudy (jdoe0001):

well... there are 1.609 km in 1 mile so we know he rides at 18mi/hr so to convert that to km we multiply it by the conversion fraction but yes, basically yes

OpenStudy (anonymous):

so how do I change what I have to make it right?

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

hold the mayo.. the 2nd needs to be a distance only

OpenStudy (jdoe0001):

\(\bf \cfrac{18\cancel{ mi}}{hr}\cdot \cfrac{1.6km}{\cancel{ mi}}\implies \cfrac{28.8km}{hr}\impliedby \textit{in 1 hr} \\ \quad \\ 4hr\cdot \cfrac{28.8km}{hr}\implies \cfrac{4\cancel{ hr}\cdot 28.8km}{\cancel{ hr}}\implies 115.2km\impliedby \textit{in 4 hrs}\)

OpenStudy (anonymous):

on the speed part the answer is set up as a blank spot you fill in and next to it is km/h so 28.8 is the speed?

OpenStudy (jdoe0001):

yeap, 28.8 is the speed, thus uses km/h unit the 4hrs is how long it has travelled, thus is distance only, thus km

OpenStudy (anonymous):

thank you so much your a lifesaver you really helped me understand this stuff!!

OpenStudy (jdoe0001):

yw